# FE-LWS: Refined Image-Text Representations via Decoder Stacking and Fused Encodings for Remote Sensing Image Captioning

### v2

@@ IV-B Numerical Evaluation
FE-LWS 0.8349 0.7589 0.6952 0.6369 0.4554 0.7767 2.9230
TABLE XI: Experimental Results of Baseline Methods and Different Searching Techniques on UCM Dataset
@@
FE-LWS 0.8805 0.8305 0.7862 0.7511 0.5014 0.8497 3.6806
TABLE XII: Experimental Results of Baseline Methods and Different Searching Techniques on RSICD Dataset
@@
Tables XII, XII and XII compares various baseline methods. R-BOW and L-FV [22] use handcrafted sentence representations with RNN and LSTM decoders, respectively. CSMLF and its fine-tuned variant [27] extend this for multi-sentence captioning. SVM-DBOW and SVM-DCONC [6] apply SVMs with bag-of-words and concatenated sentence features. TrTr-CMR [28] combines a Swin Transformer encoder with a cross-modal transformer decoder. TextGCN [3] aligns the embeddings of the decoder with frozen TextGCN output. Our model, FE-LWS, combines dual CNN encoders with stacked GRUs and local weighted aggregation to enhance semantic representation. This method improves performance by effectively integrating the encoding and decoding strategies. The results demonstrate that the proposed FE-LWS model achieves substantially superior performance compared to the baseline methods.
The qualitative evaluation is provided in the supplementary material, where the captions are evaluated by a human annotator [^1] The annotator has several years of experience in RSIC..
@@
IV-C Visual Interpretation
We present qualitative results from the test sets of three datasets in Figure 3, comparing the proposed model (FE-LWS) with ground-truth captions (Gold) and conventional encoder-decoder models (ConvNext and ResNet). In the first two examples (Figures 3a and 3b), the proposed model correctly classifies the images, while ConvNext and ResNet fail. In Figure 3a, ConvNext and ResNet misclassify the image as industrial area, whereas the proposed model accurately identifies it. In Figure 3b, ConvNext misclassifies the image as resort and ResNet as school, while the proposed model generates a meaningful and accurate caption. In the next two examples (Figures 3c and 3d), the proposed model enhances the quality of the captions by incorporating more detailed information. In Figure 3c, all models detect bridge and river, while the proposed model also captures trees. In Figure 3d, while all models detect park, the proposed model further identifies lake and road.
V Conclusion
Existing RSIC methods focus primarily on decoder enhancements, while giving limited attention to optimizing feature extraction during encoding. Moreover, most approaches rely on a single encoder and decoder, which restricts the model’s ability to handle the complex spatial and semantic characteristics of remote sensing images. To address these limitations, we propose a framework that combines features from two complementary CNN encoders and employs a stacked GRU-based decoder with a locally weighted aggregation strategy. This combination enables a more effective integration of multilevel semantic information, resulting in captions that are contextually rich and semantically accurate. The experimental results across multiple datasets demonstrate the superior performance of the proposed method compared to existing baselines. Fusion-based encoding improves spatial detail extraction, while weighted decoder stacking improves semantic coherence. Although the model shows robust performance, challenges may still arise when dealing with previously unseen or underrepresented image classes. Future research will explore directions to further enhance the generalization and robustness of the model without compromising efficiency. These advances are expected to improve the adaptability and reliability of RSIC in broader applications.
VI Datasets Used
The specifications of the three datasets used in our model are outlined below.
  • SYDNEY: SYDNEY [16] dataset consists of $613$ images, with 497 allocated for training, $58$ for testing, and $58$ for validation. It is derived from the Sydney dataset [29] and has been carefully curated to include seven distinct categories: airportindustrialmeadowoceanresidentialriver, and runway, through thoughtful cropping and selection.
  • UCM: UCM [16] dataset consists of $2,100$ images, with $1,680$ allocated for training, $210$ for testing, and 212121210 for validation. It is derived from the UC Merced Land Use dataset [30], which features $21$ land use image classes, each containing $100$ images. These classes include agricultureairportbaseball diamondbeachbuildingschaparraldenseresidentialforestfreewaygolfcourseharbourintersectionmediumresidentialmobilehomeparkoverpassparkingriverrunwaysparseresidentialstoragetanks and tenniscourt.
  • RSICD: RSICD [22] dataset includes a substantial collection of $10921$ images, with $8034$ designated for training, $1093$ for testing and $1094$ for validation. Sourced from a variety of platforms such as Google Earth [31], Baidu Map, MapABC, and Tianditu, this dataset encompasses $31$ distinct image classes, including airportbarelandbaseballfieldbeachbridgecenterchurchcommercialdenseresidentialdesertfarmlandforestindustrialmeadowmediumresidentialmountainparkparkingplaygroundpondportrailwaystationresortriverschoolsparseresidentialsquarestadiumstoragetanks, and viaduct.
VII Performance Metrices
The evaluation metrics used to assess our model’s performance are described below.
  • BLEU: Bilingual Evaluation Understudy [23] measures n-gram overlap between generated and reference captions. Calculate the geometric mean of the precision of n grams (BLEU-1 to BLEU-4) and apply a brevity penalty to discourage short outputs.
  • METEOR: Metric for Evaluation of Translation with Explicit Ordering [24] incorporates stemming, synonymy, and word order. It balances precision and recall with an emphasis on alignment and content accuracy.
  • ROUGE: Recall-Oriented Understudy for Gisting Evaluation [25] is a recall-based metric that measures n-gram overlap. We use ROUGE-L, which focuses on the longest common subsequence (LCS).
  • CIDEr: Consensus-based Image Description Evaluation [26] evaluates the similarity between generated and human-written captions by measuring consensus across multiple references, highlighting the relevance and agreement of the content.
VIII Subjective Evaluation
TABLE XIII: Subjective Evaluation of Different CNNs on Three Datasets (in %)
@@ VIII Subjective Evaluation
R W U R W U R W U
@@
FE-LWS 94.83 3.45 1.72 92.86 5.24 1.90 83.72 6.95 9.33
Unlike classification, multiple captions can describe the same image with similar effectiveness. Hence, relying solely on the test captions (five for all three data sets) to evaluate the predicted captions is not justified [22, 4]. For this reason, we performed a subjective evaluation. In this process, we assigned a human annotator with expertise in this field to assess the quality of the captions generated. This evaluation provides insights into caption quality from the perspective of human psychology. In our work, we used three labels to assess the quality of the generated captions, as discussed below.
  • Related: The generated caption accurately describes the input image.
  • Partially Related: The generated caption describes the input image but contains major errors or flaws.
  • Unrelated: The generated caption is completely incorrect and does not describe the input image in any meaningful way.
Table XIII presents the subjective evaluation results, comparing our model (FE-LWS) with conventional encoder-decoder models, where ConvNext and ResNet are used as encoders, respectively. The results indicate that our model outperforms other models.

### v3

@@ 4.2 Performance metrics
MsEdF 0.8349 0.7589 0.6952 0.6369 0.4554 0.7767 2.9230
To assess the performance of our model, we use widely used evaluation metrics in RSIC tasks, including BLEU [17], METEOR [11], ROUGE [13], and CIDEr [22].
  • BLEU: The Bilingual Evaluation Understudy (BLEU) metric [17] assesses the quality of the generated text by measuring the n-gram overlap between a generated caption and its reference captions. It calculates the geometric mean of n-gram precision scores and applies a brevity penalty to discourage excessively short outputs. BLEU is widely used in tasks such as machine translation and image captioning. In this work, we evaluate the performance using BLEU-1 through BLEU-4.
  • METEOR: The Metric for Evaluation of Translation with Explicit Ordering (METEOR) [11] is an evaluation metric that accounts for stemming, synonymy and word order to assess the similarity between generated and reference captions. Unlike BLEU, which is based purely on precision, METEOR incorporates both precision and recall, with a stronger emphasis on accuracy.
  • ROUGE: ROUGE [13] stands for Recall-Oriented Understudy for Gisting Evaluation. It is a recall-based metric that measures the overlap in n-grams between a generated caption and reference captions. In this work, we use ROUGE-L, which is based on the longest common subsequence (LCS).
  • CIDEr: CIDEr [22] stands for Consensus-based Image Description Evaluation. It is a metric used to evaluate the quality of generated captions for images by comparing them with a set of human-written reference captions. CIDEr measures how well a generated caption aligns with the collective understanding of the image across multiple human descriptions, aiming to capture the similarity between a generated caption and the general human perspective on the image content.
5 Experimental Results
This section presents a comprehensive evaluation of the proposed model, including quantitative comparisons with baseline methods and ablation studies, followed by qualitative analysis through test image visualizations and subjective evaluation.
5.1 Numerical Evaluation
Table 8: Experimental Results of Several RSIC Methods on the UCM Dataset
@@ 5.1 Numerical Evaluation
MsEdF 0.8805 0.8305 0.7862 0.7511 0.5014 0.8497 3.6806
Table 9: Experimental Results of Several RSIC Methods on the RSICD Dataset
@@
MsEdF 0.6796 0.5124 0.4073 0.3365 0.2944 0.5248 0.9423
Tables˜7, 8 and 9 compares various baseline methods. R-BOW and L-FV [16] use handcrafted sentence representations with RNN and LSTM decoders, respectively. CSMLF and its fine-tuned variant [23] extend this for multi-sentence captioning. SVM-DBOW and SVM-DCONC [7] apply SVMs with bag-of-words and concatenated sentence features. TrTr-CMR [25] combines a Swin Transformer encoder with a cross-modal transformer decoder. TextGCN [4] aligns the embeddings of the decoder with the frozen TextGCN output. Our model, MsEdF, combines dual CNN encoders with stacked GRUs and local weighted aggregation to enhance semantic representation. This method improves performance by effectively integrating the encoding and decoding strategies. The results demonstrate that the proposed MsEdF model achieves substantially superior performance compared to the baseline methods.
5.2 Ablation Studies
Table 10: Performance Comparison of Model Variants Through Ablation Studies on the SYDNEY Dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
C-NS 0.7531 0.6645 0.5982 0.5376 0.3750 0.6951 2.3954
R-NS 0.7363 0.6502 0.5785 0.5138 0.3621 0.6809 2.2526
FE-NS 0.7679 0.6814 0.6084 0.5458 0.3801 0.7263 2.4515
C-LWS 0.8141 0.7321 0.6610 0.6054 0.4369 0.7584 2.7472
R-LWS 0.8056 0.7211 0.6527 0.5924 0.4216 0.7364 2.5743
FE-LWS 0.8349 0.7589 0.6952 0.6369 0.4554 0.7767 2.9230
Table 11: Performance Comparison of Model Variants Through Ablation Studies on the UCM dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
C-NS 0.8369 0.7804 0.7312 0.6853 0.4689 0.8037 3.3990
R-NS 0.8297 0.7705 0.7178 0.6683 0.4506 0.7903 3.3329
FE-NS 0.8457 0.7919 0.7533 0.7166 0.4833 0.8163 3.4255
C-LWS 0.8655 0.8121 0.7644 0.7192 0.5086 0.8367 3.4958
R-LWS 0.8614 0.8034 0.7526 0.7049 0.4910 0.8194 3.4621
FE-LWS 0.8805 0.8305 0.7862 0.7511 0.5014 0.8497 3.6806
Table 12: Performance Comparison of Model Variants Through Ablation Studies on the RSICD dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
C-NS 0.6327 0.4636 0.3602 0.2903 0.2487 0.4648 0.8272
R-NS 0.6322 0.4621 0.3557 0.2835 0.2489 0.4649 0.8206
FE-NS 0.6452 0.4767 0.3732 0.3087 0.2609 0.4827 0.8594
C-LWS 0.6671 0.5008 0.3996 0.3302 0.2843 0.5066 0.8954
R-LWS 0.6645 0.4981 0.3963 0.3257 0.2849 0.5010 0.8684
@@ 5.2 Ablation Studies
To evaluate the contribution of individual components, we compare four encoder-decoder configurations: single encoder with single decoder, fusion-based encoder with single decoder, single encoder with LWS decoder and the proposed fusion-based encoder with LWS decoder. The results presented in Tables˜10, 11 and 12 follow the naming convention: C $\rightarrow$ ConvNext, R $\rightarrow$ ResNet and FE $\rightarrow$ Fusion of C and R. For decoder configurations, NS $\rightarrow$ a single decoder (non-stacked), while LWS $\rightarrow$ Local Weighted Stacking using three GRUs (see Section˜3.2 for details), FE-LWS $\rightarrow$ the proposed method MsEdF. The results indicate that both encoder fusion and decoder stacking improve performance, with the MsEdF configuration achieving the highest scores across all datasets. In particular, LWS offers greater performance gains than encoder fusion alone, highlighting its effectiveness even when used with a single CNN.
5.3 Visual Interpretation
@@
We present qualitative results from test images in Figure˜3. Here we have compared the proposed model (MsEdF) with ground-truth captions (Gold) and conventional encoder-decoder models (ConvNext and ResNet). These examples highlight the cases where the proposed approach produces captions that are semantically rich. Several examples (Figures˜3a, 3b and 3c) show cases where the baseline model produces incorrect or misleading captions, whereas the proposed model generates accurate and contextually relevant descriptions. In Figure˜3a, ConvNext and ResNet misclassify the image as industrial area. In Figure˜3b, ConvNext misclassifies the image as resort and ResNet as school. In Figure˜3c, ConvNext and ResNet misclassify the image as mobile home park. Other example (Figures˜3d, 3e and 3f) highlights that, although both models produce generally correct captions, the proposed approach incorporates extra visual or semantic details that enhance the overall quality of the caption. In Figure˜3d, while all models identify mountain, the proposed model further recognizes snow. In Figure˜3e, while all models identify park, the proposed model further recognizes lake and road, enhancing the semantic richness of the caption. In Figure˜3e, while all models identify ocean or river  [^1] Visually, it is challenging to distinguish whether it is an ocean or an river; therefore, both interpretations are accepted. , the proposed model also recognizes runway with marking lines.
5.4 Subjective Evaluation
Unlike classification tasks, image captioning allows multiple valid descriptions of the same image. Therefore, evaluating the generated captions solely based on the five reference captions provided in each of the three datasets is insufficient [3, 16]. To address this, we conducted a qualitative evaluation using a human annotator. [^2] A domain expert with several years of experience in RSIC. This human-centric assessment provides a deeper understanding of the quality of captions. In our study, we used three evaluation labels: Related, where the caption accurately describes the input image; Partially Related, where the caption has significant flaws but still relates to the image; and Unrelated, where the caption does not describe the image meaningfully.
Table 13: Subjective Evaluation of Different CNNs on Three Datasets (in %)
@@ 5.4 Subjective Evaluation
Related Partially Related Unrelated Related Partially Related Unrelated Related Partially Related Unrelated
@@
MsEdF 94.83 3.45 1.72 92.86 5.24 1.90 83.72 6.95 9.33
The results of the subjective evaluation of different models are presented in Table˜13. Here we compare our model (MsEdF) with conventional encoder and decoder models using ConvNext and ResNet encoders. The values indicate the percentage of test images labeled Related, Partially Related, or Unrelated based on human judgment. The results show that our model outperforms the baseline models in terms of perceptual accuracy.
6 Conclusion
Remote sensing images are challenging to interpret in the form of captions due to complex spatial layouts and ambiguous semantic objects. In this work, we propose a multi-stream encoder-decoder framework for RSIC. Two structurally diverse CNN encoders capture the varied visual characteristics present in remote sensing images. On the decoder side, a stacked GRU-based decoder is used to capture hierarchical linguistic patterns and improve the modeling of long-term dependencies. Furthermore, we show that locally weighted layer aggregation enhances semantic modeling on the decoder and enables more context-aware captioning. Experimental results demonstrate that the proposed framework consistently improves caption quality over single-stream architectures by promoting feature diversity through encoder fusion and enhancing semantic modeling via decoder-side aggregation. However, our system may fail for images that are substantially different from those encountered during training. In future work, our aim is to improve our model for unseen images through retrieval-guided decoding or domain-adaptive training.